The expected value of a normal distribution of prices for a stock is $65. If you are 95 percent sure that the price of the stock will be between $53 and $77, then what is the variance of the stock price? (Round answer to 3 decimal places, e.g. 52.750.) Variance of stock price
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut E AutoSum ー E ゴWrap Text aCopy B า 프 . Ej-., Δ. : r_一 逻锂函Merge & Center. $, % , 弼,8 conditional Format . Cell Insert Delete Format Sort &Find & 2 ClearFe Select Edting Format Painter Formatting, as Table w styles. ▼ ㆆ ▼ Clipboard Alignment Number Cells 1 100 91 92 93 94 95 96 97 98 FORMULA FOR CALCULATING THE RANGE EXPECTED VALUE -/+ (Z VALUE) (SD) AT 95% CONFIDENCE LEVEL, Z VALUE = 1.96 LOWER VALUE OF THE RANGEEXPECTED VALUE (Z VALUE)(SD) 53 = 1.96(SD) SD- 65 -1.96(SD) 12 6.12245 100 101 102 OR 103 104 105 106 107 108 109 VARIANCE-37.484 (6.12245A2) WE WILL GET SAME RESULT IF WE USE UPPER VALUE OF RANGE UPPER VALUE OF THE RANGEEXPECTED VALUE (Z VALUE)(SD) 77= 1.96(SD) SD = 65 +1.96(SD) 12 6.12245 VARIANCE-37.484 (6.12245A2) DEAR Sheet2 TIME SERIES COT REGRESSIONCAMERAEXP RETURN TREND freq MATRX INTERVAL, NORMAL HYPOTHESIS 07:26 31-01-2019